Required information (The following information applies to the questions displayed below.) In 2021, the Westgate ConstructionPLEASE HELP WITH ACCTG HW
Answer #1

Solution

2021 2022 2023
Revenue $             2,900,000 $        4,200,000 $            2,900,000
Gross profit $                696,000 $        1,008,000 $                475,600

Working

Working 2021 2022 2023
A Contract Price $ 10,000,000 $ 10,000,000 $ 10,000,000
B Cost Incurred to Date $ 2,204,000 $ 5,396,000 $ 7,820,400
C Estimated cost yet to be incurred to complete the contract $ 5,396,000 $ 2,204,000 $ 0
D = B+C Total Cost $ 7,600,000 $ 7,600,000 $ 7,820,400
E = (D/B) x 100 % of Completion 29% 71% 100%
F = A x E Revenue to date $ 2,900,000 $ 7,100,000 $ 10,000,000
G Revenue of Previous year $ 0 $ 2,900,000 $ 7,100,000
H = F - G Net Revenue this year $ 2,900,000 $ 4,200,000 $ 2,900,000
I Cost to date $ 2,204,000 $ 5,396,000 $ 7,820,400
J Cost to date of previous year $ 0 $ 2,204,000 $ 5,396,000
K = I - J Net Cost for the year $ 2,204,000 $ 3,192,000 $ 2,424,400
L = H - K Gross Profits $ 696,000 $ 1,008,000 $ 475,600